Circular No. 6815

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1998D IN NGC 5440 Y. L. Qiu, Q. Y. Qiao, W. D. Li, and J. Y. Hu, Beijing Astronom-

ical Observatory (BAO) report their discovery of an apparent supernova (mag 15.6 and 15.5 on Jan. 28.88 and 29.89 UT, respectively) on unfiltered CCD images taken with a 0.6-m telescope. SN 1998D is located at α = 14^{h}02^{m}59^{s}.28, δ = +34^{o}44′54′′.3 (equinox 2000.0), which is 26′′.1 west and 7′′.4 south of the center of NGC 5440. Unfiltered CCD images of the same field taken on Jan. 5 show no star at this position (limiting mag ∼ 18.5). A low-dispersion spectrum taken on Jan. 29.74 by J. Y. Wei (BAO) and Qiu with the BAO 2.16-m telescope at Xinglong Station shows this to be a type-Ia supernova at maximum. The spectrum displays a strong absorption feature at 615 nm. The spectrum is similar to that of SN 1994D, two days after maximum. The maximum of the Ca II triplet is at 867.6 nm.

MXB 1730–335 R. Guerriero, D. Fox and W. Lewin, Massachusetts Institute of Tech-

nology; R. Rutledge, University of California at Berkeley; C. Moore, Kapteyn Institute; M. van der Klis, Astronomical Institute ‘Anton Pannekoek’, University of Amsterdam (UA); and J. van Paradijs, University of Alabama in Huntsville and UA, report: “The All-Sky Monitor of the Rossi X-ray Timing Explorer indicates the beginning of a new x-ray outburst of this rapid burster. Discovery of a likely radio counterpart (IAUC 6813) makes timely radio and infrared observations extremely valuable.”

GS 1354–64 M. Buxton, S. Vennes, L. Ferrario, and D. T. Wickramasinghe, Aus-

tralian National University (ANU), write: “Spectroscopic observations of this transient (IAUC 6790), obtained on Jan. 20, 22, and 23 at the ANU 2.3-m telescope, reveal the emission of strong Hα, Hβ and He II 468.6-nm and weaker He I 447.1-nm and Hγ lines. The continuum is featureless and heavily reddened, and the flux at 700.0 nm corresponds to R = 17.4 ± 0.3. The Hα profile varied from double- to single-peaked over the three nights of observation, reminiscent of the behavior of the blackhole candidate GRO 1655–40 during its 1996 outburst and strongly suggestive of an origin in an accretion disk. The separation between the peaks of Hα (Jan. 20) was measured to be 600 km/s. We possibly detected a narrow red-shifted Hα absorption line (perhaps from the companion star) adjacent to the emission feature. The period of GS 1354–64 may be on the order of days.”

1998 January 30  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green